Help me solv this?
Jennifer is six years older than Sue. In four years she will be twice as old as Sue was five years ago. Find their ages now.

Answer by ptaylor(2198)

Let x=Sue's age
then x+6=Jennifer's age
x+10=Jennifer's age in 4 years
x-5 =Sue's age 5 years ago
So our equation to solve is:
x+10=2(x-5)
x+10=2x-10
x=20-------Sue's age
x+6=26 Jennifer's age
CK
26+4=2*(20-5)
30=30
Hope this helps---ptaylor